Grade 10Trigonometry

sin cos tan 75?

balsingh bassi
13 Years agoGrade 10

2 Answers

Ganga sagar Kothashiv
13 Years ago

(31/2+1)/(2*21/2)

(31/2-1)/(2*21/2)

(31/2+1)/(31/2-1)

Aman  Bansal
13 Years ago

Dear Balsingh,

sin(75)=sin(45+30)
=sin45cos30+sin30cos45
=sqrt(2)/2*sqrt(3)/2+1/2*sqrt(2)/2
factor out a sqrt(2) and the 1/4 and you get
Sin75° = √2/4(1+√3)

cos(75)=cos(45+30)
=cos45cos30-sin45sin30
=sqrt(2)/2*sqrt(3)/2-sqrt(2)/2*1/2
factor out a sqrt(2) and the 1/4 and you get
Cos75° = √2/4(√3-1)

tan75=sin75/cos75
=√2/4(1+√3)/√2/4(√3-1)
the √2/4 goes away
so it''s (1+√3)/(√3-1)
rationalize the denominator by multiplying (1+√3) on the top and bottom
you get (1+√3)^2/2 after simplifying
=(1+2√3+3)/2
=√3 +2

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert